HUBBLE SPACE TELESCOPE DAILY REPORT #3097 PERIOD COVERED: 0000Z (UTC) 04/17/02 - 0000Z (UTC) 04/18/02 Daily Status Report as of 108/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: The following scheduled proposals did not execute properly (the take data flag remained down) due to the Gyro #3 problem described in yesterday's report: WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) STIS/CCD 8904 (Bias Monitor-Part 2) Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) Five Sets ACS/WFC 9018 (Flat Field Stability) STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) STISMA1/MA2 8965 (MAMA Darks) STIS/CCD 8906 (Hot Pixel Annealing) STIS/MA1 8922 (FUV-MAMA Cycle 10 Flats) ********************************************************** The following proposals partially executed once FGS control was achieved: 1.1 WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. Fine lock was achieved approximately three hours into this 15-hour proposal. 1.2 ACS/WFC 9018 (Flat Field Stability) The Advanced Camera for Surveys (WFC) was used to assess the stability and uniformity of the low-frequency flat fields {L-flat} of the WFC detector by using multiple pointing observations of the globular clusters 47 Tucanae {NGC104}, thus imaging moderately dense stellar fields. By placing the same star over different portions of the detectors and measuring relative changes in its brightness it will be possible to determine local variations in the response of the detectors. Fine lock was achieved approximately two hours into this three-and-one-half-hour proposal. ********************************************************** 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 1 Scheduled Re-acquisitions: 9 Successful: 2 Following the Gyro #3 problems, Gyro #1 was powered on at 107/1013Z. The gyro configuration was changed at 107/1349Z to reflect active gyros #3, #4, #2, #1 with gyro #3 in the shadow mode. Fine attitude control was established via an attitude reference update at 107/1710Z with the first successful re-acquisition at 107/1713Z. A subsequent re-acquisition at 107/1750Z was also successful. Then, at 108/0749Z, there was a successful acquisition. 2.2 FHST Updates: Scheduled: 3 Successful: 3 2.3 Operations Notes: See 2.1 for events relating to the recovery from the Gyro #3 problem. The 486 engineering status buffer limit for SESBSLD was updated eleven times per ROP DF-18A. Using ROP NS-3, the NSSC-1 status buffer was dumped at 107/1009Z. The Battery #3 Capacity Test began at 107/1124Z and is proceeding nominally. Per HSTAR 8614, at 107/171406Z, during the first successful re-acquisition, a measurement of the V1 (roll) error while guiding was made. This measured value (QSFGSENB0) was -2188.939 a-s was saved, since the error null bias is not calculated for a re-acquisition. At 107/173750Z, QDVTHEJ1 (Observer Rate Bias Estimate 2) flagged out-of-limits with a value of -3.70619e-10 and QDVTHEJ2 (Observer Rate Bias Estimate 3) flagged out-of-limits with a value of 1.1934e-9. HSTAR 8613 was written. The following operations requests were completed during the day: 16671-0 - ARU/PRT @ 107/0904z 16672-1 - Gyro 1 Power On @ 107/1013z 16674-0 - ARU/PRT @ 107/1233z 16673-0 - Gyro 3 Reconfiguration @ 107/1349z 16667-0 - HLGBU from LBBias Maps @ 107/1710z 16675-0 - R/T Map @ 107/1755z 16678-0 - ARU/PRT @ 107/1913z 16679-0 - R/T Map @ 107/1929z 16681-0 - R/T Map @ 107/2111z 16624-0 - Reset NCS Min/Max buffer @ 107/2112z 16682-0 - ARU/PRT @ 107/2237z 16684-0 - ARU/PRT @ 108/0109z 16685-0 - HLGBU from LBBias Maps @ 108/0121z 16686-0 - ARU/PRT @ 108/0216z 16687-0 - R/T Map @ 107/0345z 16691-0 - ARU/PRT @ 108/0435z 16689-0 - R/T Map @ 108/0513z 16680-1 - Mask -E SPA from TRSWCC stack @ 108/0523z 16662-0 - Battery 3 Reconditioning (thru step 30) @ 108/0524z 16690-2 - HLGBU from LBBias Maps @ 108/0528z 16666-0 - Recorder Management for Battery Reconditioning (Completed Continuous Engineering Record Mode) @ 108/0538z 16692-0 - HLGBU from LBBias Maps @ 108/0713z 16693-0 - ARU/PRT @ 108/0714z The following operations notes were completed: 989-0 - Gyro 3 Reconfiguration Limits 990-1 - Reset Safing Limits for Gyro Reconfiguration 991-0 - Ignore ACS STB 290 due to Event Flag2 being set 911-0 - Limit Management During WFPC2 Decontamination (M001) 3.0 SIGNIFICANT FORTHCOMING EVENTS: The NICMOS recovery-is scheduled to begin tomorrow at approximately 109/0705Z. Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW